OpenStudy (anonymous):

Find the critical numbers of the function f(x) = x^6(x − 1)^5

OpenStudy (anonymous):

... requires implicit differentiation...

OpenStudy (jmark):

Given function f(x) = x^6(x − 1)^5 Take the derivative: f'(x) = 6x^5(x - 1)^5 + 5x^6(x - 1)^4 Set the derivative equal to 0 and solve: 6x^5(x - 1)^5 + 5x^6(x - 1)^4 = 0 (x^5(x - 1)^4) [6(x - 1) + 5x] = 0 (x^5(x - 1)^4) (x - 6) = 0 x = 0 and x = 1 and x = 6 Hence these are the critical points of the function ICSE 9th Class Question Papers - http://scoremore.beep.com/-icse-9th-class-question-papers-2014-07-11-1.htm